### v2

@@ Abstract
Service robots work in a changing environment habited by exogenous agents like humans. In the service robotics domain, lots of uncertainties result from exogenous actions and inaccurate localisation of objects and the robot itself. This makes the robot task scheduling problem challenging. In this article, we propose a benchmarking framework for systematically assessing the performance of algorithms scheduling robot tasks. The robot environment incorporates a map of the room, furniture, transportable objects, and moving humans. The framework defines interfaces for the algorithms, tasks to be executed, and evaluation methods. The system consists of several tools, easing testing scenario generation for training AI-based scheduling algorithms and statistical testing. For benchmarking purposes, a set of scenarios is chosen, and the performance of several scheduling algorithms is assessed. The system source is published to serve the community for tuning and comparable assessment of robot task scheduling algorithms for service robots. The framework is validated by assessment of scheduling algorithms for the mobile robot executing patrol, human fall assistance and simplified pick and place tasks.
@@
[…]
{graphicalabstract}[Uncaptioned image]
{highlights}
Goal: Comparable evaluation of various algorithms scheduling diverse robot tasks.
Method: Kinematic simulation executed in configured or generated scenarios and environments.
Result: The framework teaches an RL-based algorithm, and other algorithms are compared.
Conclusion: The comparison shows algorithm performance for specific scenarios and tasks.
1 Introduction
@@ 1.1 Problem statement
Humans have honed the ability to manage and prioritise multiple tasks for centuries. Today, the challenge lies in transferring this capability to robots. This task has become increasingly important due to the growing need for robotisation in various domains facing staff shortages. Thus, robots ability to multitask efficiently becomes critical, and the proposed scheduling algorithms should be assessed systematically.
Solving the problem of robot task management requires adapting task scheduling solutions to real-world environments characterised by dynamic changes [8], uncertainty [13], and interactions with exogenous actors such as humans [12, 18]. This presents a significant challenge, as changing environments affect the schedules and action plans required to perform the requested tasks. For safety and effectiveness reasons, task scheduling algorithms should be rigorously tested in simulated environments [3, 16]. Evaluations of these algorithms may involve statistical analysis or proving their optimality. The latter is especially difficult in changing, inhabited environments, particularly for neural-network-based scheduling algorithms [15].
Task scheduling for mobile robots in dynamic environments has been extensively studied, with researchers exploring both traditional and AI-based methods to address the inherent challenges. Classical approaches often rely on heuristic or optimization-based strategies. However, these methods struggle with scalability and adaptability in environments with frequent changes. Modern AI-driven methods, such as reinforcement learning, have shown promise in addressing dynamic scheduling. For example, [30] review studies focused on Reinforcement Learning techniques that have been used for dynamic task scheduling. [33] argue that a Hybrid Q-learning approach allows for maintaining optimal scheduling. Furthermore, they state their reinforcement learning method is resilient to non-deterministic environments. They validate the algorithm in a robotic arm environment, and in the future, they want to focus on motion planning, collision avoidance, and multi-agent algorithms. Similarly, [36] proposed a deep multi-agent reinforcement learning approach to optimize task scheduling in human-robot collaboration, showcasing its effectiveness in dynamic environments. Yet, these techniques often require extensive training data and may falter in highly unpredictable scenarios. Simulation tools like Gazebo and frameworks such as ROSPlan [5] have facilitated the testing of scheduling algorithms, but they are limited in their ability to capture the complexities of human-robot interactions and real-world uncertainties. Despite these advancements, benchmarking service robot scheduling algorithms in such environments remains an open problem, as existing tools and methods fall short in addressing the unpredictable and stochastic nature of real-world settings, and transparent comparison of the methods is difficult. Thus, a framework for comparable benchmarking of the algorithms is mandatory.
Although current frameworks and benchmarking systems provide essential tools for evaluating task scheduling algorithms, they often fall short in addressing the complexities of dynamic and uncertain environments in the field of service robotics. The framework introduced in this work aims to bridge this gap by offering a structured approach to training and benchmarking scheduling algorithms. The proposed framework explicitly accounts for real-world uncertainties and facilitates fair performance comparisons across diverse scenarios. The framework is adaptable, supports various environments and algorithms, and allows for systematic and repeatable testing against configurable scenarios. Importantly, the framework ensures that performance scores are comparable between different algorithms and scenarios. In addition, it provides tools for generating test cases and scenarios, which can be used to train AI-based scheduling algorithms. Additionally, this article introduces a SysML-based [35] domain-specific language (DSL) designed to support the presentation, documentation, and analysis of benchmarking systems developed with the proposed system. The DSL follows the Model-based Systems Engineering (MBSE) method used broadly for systems analysis, design verification and validation in the V-model [6] workflow. The framework’s capabilities are illustrated through the use case diagram Fig. 1. As target users of the framework, we recognise the scientists developing scheduling algorithms and mobile manipulator integrators. The algorithm developers can use the framework to test their solutions against other algorithms and previous versions of their work in a comparable environment. The mobile manipulator integrators can mimic a use-case-specific environment in the framework and validate which scheduling algorithms best suit the tasks and environment of the robot application.
@@
The framework should ease analysing how the scenario’s change influences the algorithm’s score. This article investigates how such a system should be structured, how it works, and how to use it to receive comparable results of different scheduling algorithms evaluated in various scenarios.
The best-assessed scheduling algorithm must be integrated into the end product— robot control system. For the architecture-level integration, we have proposed the Simulation-Physical Modeling Language [9], and for the implementation-level integration, we have published the ROS [^1] Robot Operating System [19, 34]-based TaskER framework [10] that models the structure and behaviour of a multitasking robot. TaskER manages safe interruptions between different task contexts, allowing the robot to conduct safety-critical actions of the interrupted task before executing the switch to the next task.
The contribution of this article can be summarized as follows:
  • a comprehensive framework for training and benchmarking robot task scheduling algorithms, explicitly addressing real-world uncertainties in dynamic and human-inhabited environments. It supports systematic, repeatable testing across configurable scenarios while ensuring comparable performance evaluation.
  • a dedicated domain-specific language for presenting, documenting, and analyzing benchmarking systems, enhancing transparency and standardization in the development and assessment of scheduling algorithms.
  • an Open-Source and Practical Validation: The framework, validated through scenarios involving patrol, human fall assistance, and pick-and-place tasks, is published as an open-source tool, promoting community collaboration and the development of AI-based scheduling algorithms.
In the next section 2, we summarise the state-of-the-art solutions to the problem of benchmarking task scheduling algorithms. In the following sections, we introduce the architecture of the proposed benchmarking system 3.1, the scope of possible benchmarking scenarios 3.2, the configuration of the system for testing specific algorithms and robots 3.4, and finally, we describe the developed tools for generating test-cases 3.5. In section 4, we describe the configuration of the example system that evaluates basic and neural-network-based algorithms. The neural network algorithms were taught on the test cases generated by the proposed tools. Finally, the evaluation results are presented and discussed in 4.2.
2 Related work
2.1 Simulators for developing and benchmarking in robotics
Simulators play a crucial role in developing and benchmarking algorithms in the robotics domain. The Stage simulator, as discussed by [14], provides a simplified platform for benchmarking navigation systems in static environments. However, its project has been inactive for a long time, making its utility limited by the challenges of integrating recent algorithms. Additionally, the review of physics-based simulators by [23] emphasizes the importance of simulation accuracy in testing robotic controllers and the potential benefits of simplifying simulations for high-level task scheduling problems. The varying performance of simulators across different robotic domains, as noted by [7], further underscores the need for tailored benchmarking systems that account for the specific physical effects relevant to each application.
Simulation is a crucial development tool for robotics, which allows avoiding challenges of learning on the real robot, especially when learning-based techniques are applied [17]. This approach avoids the time-consuming experiments on real platforms and mitigates the risk of damaging the real robots. Moreover, the utilization of parallel architecture of the Graphical Processing Units enables simulating many robots at the same time and significantly reduces the training time [28]. The new PhysX-based Orbit platform standardizes the benchmarking manipulation and robot locomotion tasks [25].
2.2 Benchmarking task scheduling
The problem of robot task scheduling in dynamic and uncertain environments has garnered significant attention in the field of robotics. In the general domain of task scheduling, several frameworks and benchmarking systems have been proposed. One notable framework is introduced by [27], which leverages large language models (LLMs) for task division and tool selection. The framework evaluates the performance of three LLMs by measuring the success rate and the order of tool usage in completing subtasks. This approach demonstrates the potential of LLMs in enhancing task scheduling through intelligent tool selection, although its primary focus is on utilising language models rather than the broader context of robot task scheduling. LLMs are also used to interpret vocal requests of tasks and identify their parameters [32]. Another significant contribution is TASKOGRAPHY [1], a large-scale benchmark designed to evaluate robotic task planning over 3D scene graphs (3DSGs). This work also introduces SCRUB, a planner-agnostic strategy for adapting 3DSGs to improve planning performance, and SEEK, a procedure that enhances learning-based planners’ ability to exploit 3DSGs. These tools are particularly relevant for environments requiring detailed scene descriptions, although they focus more on planning than scheduling tasks in a dynamic environment.
The survey by [4] provides a comprehensive overview of various task scheduling metrics, offering a foundational understanding of the criteria used to assess scheduling algorithms. This work, along with the review of metrics and benchmarking for parallel job scheduling by [11], highlights the diversity of metrics applicable to scheduling, though these works primarily concentrate on static environments. In the domain of task and motion planning (TAMP), [21] proposes a platform-independent evaluation method, which is crucial for assessing the generalizability of scheduling algorithms across different robotic platforms. Similarly, the work by [27] offers a structured framework for LLM-based AI agents, focusing on the execution of inference processes, but again, with a specific focus on language models rather than broader task scheduling in robotics.
For benchmarking dynamic job scheduling, [24] presents the Dynamic Job Scheduler Benchmark (DJSB), a tool for comparing performance metrics across different scenarios. This tool, which allows for testing various resource management strategies, is particularly relevant for applications where resource availability and task demands change over time. The framework proposed by [2] addresses the PET-aware task-to-core scheduling problem, offering evaluation procedures and benchmarks for comparing scheduling algorithms. This work, along with the graph-based scheduling algorithm benchmarking by [20], contributes valuable methodologies for assessing the efficiency and performance of scheduling algorithms, particularly regarding resource utilization and execution time.
In summary, while existing frameworks and benchmarking systems offer valuable tools and methodologies for evaluating task scheduling algorithms, there remains a need for a comprehensive system that addresses the unique challenges posed by dynamic and uncertain environments in the service robotics domain. Our proposed framework seeks to fill this gap by providing a systematic approach to training and benchmarking scheduling algorithms, incorporating real-world uncertainties and enabling fair comparisons across different scenarios. Besides the above core functional contribution, this article provides a domain-specification language (DSL) for the presentation, documentation, and analysis of the benchmarking systems that are built using the proposed framework. The DSL is expressed in Systems Modeling Language (SysML) [35].
@@ 2.2 Benchmarking task scheduling
3 The framework for benchmarking and training scheduling algorithms
3.1 The architecture
TaBSA— the framework for Training and Benchmarking Scheduling Algorithms is a metamodel that, by configuration, gives a specific benchmarking system named TaBSA System. The TaBSA System is organised in a way presented in the SysML block definition diagram in Fig. 2. The diagram is the part of the metamodel for the variety of particular TaBSA System configurations (realisations). For each block, a dedicated stereotype was introduced to simplify the specification of these realisations. The TaBSA System addresses the use cases [UCX] from Fig. 1, by, a.o., reflecting the entities from these use cases as dedicated stereotypes: <<TaBSASystem>> – [UC1] and [UC2], <<Scenario>> – [UC1.2], <<Robot>> – [UC1.2.1], <<Task>> – [UC1.2.2], <<Environment>> – [UC1.2.3], <<DecAgent>> – [UC1.3], <<EvalFunction>> – [UC1.4].
Figure 2: TaBSA system structure
@@ 3.1 The architecture
The <<TaBSASystem>> general behaviour is also a part of the metamodel. It is depicted in activity diagram 3. It provides a scheme for repeatable execution of configurable scenarios [UC1]. Activity [ACT1] addresses configurability, [ACT2] in particular addresses teaching and a step of the task execution [UC2], [ACT3] evaluates the latest task change decision [UC1.4]. A detailed description of the particular operations mentioned in the diagram is presented in the following part of the article.
Figure 3: TaBSA system main operation
3.2 Scope of benchmark scenarios
@@
The ’tasks’ list contains every <<Task>> the robot should work on during the specific training session. We can define the different types of <<Task>>s our robot is supposed to perform as long as they can be represented in our restricted environment, as per [UC1.2.2]. This includes, a.o., simple operations such as navigation, picking, and placing of objects. Once we decide, we can fill the list with the selected amount for each type of <<Task>>. Each <<Task>> must possess the ’request_time’ property that represents at what time it is presented to the robot. The <<Task>>’s operations include is_called() and is_completed(). Apart from ’tasks’, the scenario also possesses the ’jobs’ list. This list contains solely the <<Task>>s that have already been called (is_called() operation returns True), so it can never contain a <<Task>> that can’t be found on the ’tasks’ list.
The <<Scenario>> defines the ’environment’ (of <<Environment>> stereotype) consisting of a flat map with walls, door openings, and furniture (simple shapes blocking the movement), items placed on the furniture, and moving humans represented by footprints of their legs, as per use case [UC1.2.3]. Simulated humans move on the map using simple kinematic equations and PRM algorithm for path planning, and their behaviour defines what happens after they reach the end of their path. They can circle between two points, choose a new destination, instantly change their position and choose a new destination, or just disappear. Humans are dynamically embedded on the map to make it possible for the robot to consider them during navigation. This represents the system incorporating external sensors for human detection. Manipulation targets are described solely by objects’ location. These provide a highly simplified representation of a real robot’s environment but are sufficient to evaluate <<Task>> execution order and allow training of neural network-based algorithm in a reasonable time. Due to its simplicity, it also satisfies the use case [UC2.2].
Table 1: Task operations as called by different components of our example benchmarking system — Mobile Manipulator System (for example pseudocodes see Fig. 4)
Operation Called by
is_called() <<Scenario>> - to determine when to add the <<Task>> to ’jobs’ list
is_completed() <<Scenario>> - to determine when to remove <<Task>> from ’jobs’ list
DQN Eval (Fig. 7) - to determine when the <<Task>> is completed and proper reward should be granted
Statistic Eval (Fig. 7) - to determine when the <<Task>> is completed to calculate time difference between
     planned and final completion
is_alive() DQN Eval (Fig. 7) - to determine if the <<Task>> died and therefore <<Scenario>> should terminate
Statistic Eval (Fig. 7) - to determine if the <<Task>> died and therefore <<Scenario>> should terminate
work_for() Mobile Manipulator (Fig. 7) - inside it’s execute_step() operation, it is passed as argument by the
     <<Scenario>>’s work() operation
estimate_duration() <<Scenario>> - to update <<Task>>’s estimated durationbased on it’s own properties. It is called
     by update_job() operation
3.3 Task features
@@
$deathtime=deadline+max\_delay.$ (1)
The is_alive() operation returns False if the <<Task>> has died, i.e. it has exceeded its deadline by maximum delay. The <<Task>>s whose execution delay does not lead to critical outcomes can’t die. The estimate_duration() operation purpose is to determine the theoretical duration of a <<Task>> based on its properties. The duration is extended by the minimum task duration passed as an argument to the operation. Use cases for these operations are outlined in Table 1. The properties of the example <<Task>>s are as described:
@@ 3.4 Benchmark configurability
3.4.1 Configuring the Decision Agent
The <<DecAgent>> has one basic operation select_task() for selecting a <<Task>> for execution. To make this decision, it receives the list of jobs, the current time, and the output of the <<EvalFunction>> from the previous decision. The job list contains tasks already submitted but not completed (the ’jobs’ list of the <<Scenario>>). The list may contain more <<Task>>s than the <<DecAgent>> can process. If this is the case, it is at the <<DecAgent>>’s discretion to limit the number of <<Task>>s to be considered. The functionality of the agents can be extended using <<AgentPlugin>>s.
@@ 3.4.3 Creating the Agent Plugins
An <<AgentPlugin>> extends an agent with a selected capability of considerable complexity. Each <<AgentPlugin>> has its unique interface, depending on how it works. This is due to the infinite number of potential skills with which a <<DecAgent>> can be extended. <<DecAgent>>s wishing to use them must adapt to the way it works. By separating a capability as a <<AgentPlugin>>, there is no need to implement it individually for each <<DecAgent>>, and new <<DecAgent>>s are not forced to expand the code of the old ones to have it. Any <<DecAgent>> can use these <<AgentPlugin>>s. An illustrative example of an <<AgentPlugin>> could be a ’task request predictor’ that estimates the timestamps of future tasks.
@@ 3.4.5 Configuring the Robot
3.5 Test-case configuration
@@ 3.5.1 The robot’s environment generation
The <<Environment>> in which the <<Robot>> works is <<Scenario>>’s property. The static obstacles emulating walls for the test session are generated using the recursive division method configured with the provided parameters (Fig. 5). First, all rooms, walls, and doors are generated considering the provided constraints. After that, the number of pieces of furniture is added to the map so as not to stand in front of any door. The number of furniture pieces is limited but not constant, as some rooms may be too small to place that much furniture, or the furniture just won’t generate properly in a finite number of tries. A set number of manipulation-intended items is placed on each piece of furniture. Ultimately, we generate starting and destination poses for human actors and define their behaviour.
Figure 5: Configurable environment parameters
The probability map is calculated based on the pedestrian spawning probabilities set during the environment configuration (Fig. 5/Pedestrians). Different probabilities can be set for the rooms, doors, and map entrance. The example environment map is presented in Fig. 6. All of this satisfy [UC2.3], specifically [UC2.3.1] and [UC2.3.2].
Figure 6: Sample randomly generated map with objects (blue), calculated probability distribution for pedestrian spawning (the darker green the bigger probability), and example footprints of the pedestrians (red) at the system runtime

### v3

@@ Abstract
This article introduces a software framework for benchmarking robot task scheduling algorithms in dynamic and uncertain service environments. The system provides standardized interfaces, configurable scenarios with movable objects, human agents, tools for automated test generation, and performance evaluation. It supports both classical and AI-based methods, enabling repeatable, comparable assessments across diverse tasks and configurations. The framework facilitates diagnosis of algorithm behavior, identification of implementation flaws, and selection or tuning of strategies for specific applications. It includes a SysML-based domain-specific language for structured scenario modeling and integrates with the ROS-based system for runtime execution. Validated on patrol, fall assistance, and pick-and-place tasks, the open-source framework is suited for researchers and integrators developing and testing scheduling algorithms under real-world-inspired conditions.
@@
robot scheduling algorithms , benchmarking tools , dynamic environments , reinforcement learning
journal: SoftwareX
Metadata
Nr. Code metadata description Please fill in this column
C1 Current code version 1.0.0
C2 Permanent link to code/repository used for this code version https://github.com/RCPRG-ros-pkg/Smit-Sim/tree/v1.0.0
C3 Permanent link to Reproducible Capsule https://github.com/RCPRG-ros-pkg/Smit-Sim/blob/v1.0.0/Dockerfile
C4 Legal Code License MIT License
C5 Code versioning system used git
C6 Software code languages, tools, and services used python3.7, bash, ROS melodic
C7 Compilation requirements, operating environments & dependencies https://github.com/RCPRG-ros-pkg/Smit-Sim/blob/main/requirements.txt
C8 If available Link to developer documentation/manual https://github.com/RCPRG-ros-pkg/Smit-Sim/blob/main/README.md
C9 Support email for questions daniel.gieldowski@pw.edu.pl
Table 1: Code metadata (mandatory)
1 Motivation and significance
@@ 1.1 Problem statement
The growing demand for robotisation in sectors facing labour shortages highlights the need for autonomous robots capable of managing multiple tasks efficiently. A critical challenge in this context is designing and evaluating scheduling algorithms that can operate reliably in dynamic [1], uncertain [2], and human-inhabited environments [3, 4]. While AI-based approaches [5, 6], such as reinforcement learning, offer promising solutions, their real-world deployment remains difficult due to the unpredictability of operational contexts and the lack of reliable, comparable evaluation tools.
Existing simulation platforms and benchmarking frameworks, although useful, often fall short in capturing the complexity of real-world conditions and ensuring transparent comparison of algorithmic performance. As a result, the assessment of scheduling algorithms lacks consistency and reproducibility, limiting their integration into practical robotic systems.
The goal of this work is to provide a software framework that enables systematic and reproducible evaluation of both existing and emerging scheduling approaches, offering a solid foundation for the development and comparison of new algorithms. Demonstrating strong performance in some scenarios or failure in others are both valuable outcomes, as they highlight the framework’s ability to reveal algorithmic strengths and weaknesses in dynamic and uncertain environments. To facilitate adoption, we provide a Dockerfile for easy deployment and an introductory video [^1] https://vimeo.com/1122196556 .
1.2 Software capabilities
The software framework is dedicated to training and benchmarking robot scheduling algorithms. It enables systematic, repeatable testing in configurable scenarios and ensures fair performance comparisons across algorithms. The framework supports uncertainty modeling, task-specific scenario generation, and comparative evaluation, making it suitable for both algorithm developers and mobile manipulator integrators. It also includes a Domain-Specific Language (DSL) based on Systems Modeling Language (SysML), which facilitates structured documentation and analysis of benchmarking setups using a model-based systems engineering approach. Additionally, the SysML-based description supports assessing software consistency between simulation and physical deployment via SPSysML [7].
This article presents the framework’s architecture, usage, and validation. It also outlines its integration with robot control systems, including the ROS-based TaskER platform [8] for safe task switching. The proposed framework aims to standardize and accelerate the development of robust, multitasking robots for real-world applications. The framework’s capabilities are illustrated through the use case diagram Fig. 1.
@@ 1.2 Software capabilities
The proposed framework facilitates benchmarking and development of task scheduling algorithms in realistic robotic environments. It has been validated in three service robot scenarios:
  • 1.
    patrol and monitoring tasks in dynamic spaces,
  • 2.
    assistance in human fall detection and intervention,
  • 3.
    pick-and-place operations involving object manipulation.
We provide ready-to-use scenario generators, benchmarking tools, and an interface for training AI-based algorithms under controlled uncertainty. Researchers and integrators can test algorithms in reproducible settings and evaluate their robustness to changing conditions.
The main contribution of this work can be summarized as follows:
  • 1.
@@
    a complete framework for training and benchmarking robot task scheduling algorithms in uncertain, human-inhabited environments,
  • 2.
    integration of a domain-specific language for formal documentation and analysis of benchmarking systems,
  • 3.
    open-source implementation validated in real-world-inspired service robot scenarios, enabling repeatable and comparable evaluation of scheduling methods.
2 Software description
TaBSA— the framework for Training and Benchmarking Scheduling Algorithms bases on a metamodel that, by application, gives a specific benchmarking system named TaBSA System. The TaBSA System is organised in a way presented in the SysML block definition diagram in Fig. 2(a).
(a) TaBSA system structure
(b) TaBSA system main operation
Figure 2: TaBSA structure and behaviour
@@ 2 Software description
The <<TaBSASystem>> general behaviour is also a part of the metamodel. It is depicted in activity diagram 2(b). It provides a scheme for repeatable execution of configurable scenarios [UC1]. Activity [ACT1] addresses configurability, [ACT2] in particular addresses teaching and a step of the task execution [UC2], [ACT3] evaluates the latest task change decision [UC1.4]. A detailed description of the particular operations mentioned in the diagram is presented in the following part of the article.
2.1 Framework configurability
2.1.1 Scope of benchmark scenarios
@@
The ’tasks’ list contains every <<Task>> the robot should work on during the specific training session. We can define the different types of <<Task>>s our robot is supposed to perform as long as they can be represented in our restricted environment, as per [UC1.2.2].
The <<Scenario>> defines the <<Environment>> consisting of a flat map with walls, door openings, and furniture (simple shapes blocking the movement), items placed on the furniture, and moving humans represented by footprints of their legs, as per use case [UC1.2.3]. Simulated humans move on the map using simple kinematic equations and PRM algorithm for path planning, and they execute complex behaviours like going though list of destinations. Humans are embedded on the map, so the robot has to avoid them. Robot manipulation targets placed in the environment are moveable by the robot. These provide a highly simplified representation of a real robot’s environment but are sufficient to evaluate <<Task>> execution order and allow training of neural network-based algorithm in a reasonable time. Due to its simplicity, it also satisfies the use case [UC2.2].
2.2 Tasks adaptation
@@
deathtime=deadline+max_delay.deathtime=deadline+max\_delay. (1)
The properties of the example <<Task>>s are as described:
  • 1.
@@ 2.3 Benchmark configurability
2.3.1 Configuring the Decision Agents
The <<DecAgent>> has one basic operation select_task() for selecting a <<Task>> for execution. To make this decision, it receives the list of jobs, the current time, and the output of the <<EvalFunction>> from the previous decision. The job list contains tasks already submitted but not completed (the ’jobs’ list of the <<Scenario>>). The list may contain more <<Task>>s than the <<DecAgent>> can process. If this is the case, it is at the <<DecAgent>>’s discretion to limit the number of <<Task>>s to be considered. The functionality of the agents can be extended using <<AgentPlugin>>s. In one benchmark, you can execute different Decision Agents at once for comparison.
@@ 2.3.3 Creating the Agent Plugins
An <<AgentPlugin>> extends an agent with a selected capability of considerable complexity. Each <<AgentPlugin>> has its unique interface, depending on how it works. This is due to the infinite number of potential skills with which a <<DecAgent>> can be extended. By separating a capability as an <<AgentPlugin>>, there is no need to implement it individually for each <<DecAgent>>, and new <<DecAgent>>s are not forced to expand the code of the old ones to have it. Any <<DecAgent>> can use these <<AgentPlugin>>s. An illustrative example of an <<AgentPlugin>> could be a ’task request predictor’ that estimates the timestamps of future tasks.
@@ 2.3.5 Configuring the Robot
2.4 Test-case generators
The test case configuration that is the occupancy map, human trajectories, and the real— unknown to the robot, task schedule, can be set manually for a given specific application, or can be generated with the prepared generators.
@@ 2.4.1 The robot’s environment generation
The <<Environment>> in which the <<Robot>> works is <<Scenario>>’s property. The static obstacles emulating walls for the test session are generated using the recursive division method configured with the provided parameters (Fig. 4(a)). First, all rooms, walls, and doors are generated considering the provided constraints. After that, the number of pieces of furniture is added to the map so as not to stand in front of any door. The number of furniture pieces is limited but not constant, as some rooms may be too small to place that much furniture, or the furniture just won’t generate properly in a finite number of tries. A set number of manipulation-intended items is placed on each piece of furniture. Ultimately, we generate starting and destination poses for human actors and define their behaviour.
(a) Configurable environment parameters
(b) Example randomly generated map with objects (blue), calculated probability distribution for pedestrian spawning (the darker green the bigger probability), and example footprints of the pedestrians (red) at the system runtime
Figure 4: Robot’s environment configuration
The probability map is calculated based on the pedestrian spawning probabilities set during the environment configuration (Fig. 4b/Pedestrians). Different probabilities can be set for the rooms, doors, and map entrance. The example environment map is presented in Fig. 4(b). All of this satisfy [UC2.3], specifically [UC2.3.1] and [UC2.3.2].